# A vectorial Darboux transformation for integrable matrix versions of the Fokas-Lenells equation

### v1

@@ 6.2.2 The triple root case and rogue waves
Figure 10: Plots of the absolute values of the two components of a solution from the class in Example 6.7, both showing rogue waves with dark and anti-dark parts. Here we chose $\alpha_{1}=1$, $\alpha_{2}=2$, $A_{1}=1/3$, $A_{2}=\sqrt{2}/3$, $c_{1}=0$, $c_{2}=c_{3}=1$.
@@
Figure 11: Plots of the absolute values of the two components of a solution from the class in Example 6.7, exhibiting rogue wave pairs. Here we chose the same data as in Fig. 10, but now $c_{1}=1$ and $c_{3}=0$.
@@
Again, we assume that $r$ is a triple root of the cubic equation, hence given by (6.54). Then $\gamma$ and $A_{\mu}$, $\mu=1,2$, are given by (6.55) and (6.56), respectively. Now the linear system takes the form
@@
where $f=(f_{1},\ldots,f_{6})^{T}$ satisfies $f_{t}=Cf$ with the constant nilpotent matrix
@@
Inserting the expressions for $\chi_{1},\chi_{2}$ and $\Omega^{-1}$ in (6.49), determines a family of quasi-rational solutions of the two-component vector FL equation, which are localized on the plane wave background. These are thus rogue waves. They are of second order in the sense that here $\Gamma$ is a $2\times 2$ Jordan block, which causes the appearance of higher (than in the $n=1$ case) order polynomials of the dependent variables. What we obtained here has to be distinguished from a (nonlinear) superposition of two first order rogue waves from Example 6.7. $\Box$

### v2

@@ 6.2.3 The triple root case
Figure 12: Plots of the absolute values of the two components of a solution from the class in Example 6.15, both showing rogue waves. Here we chose $\alpha_{1}=1$, $\alpha_{2}=2$, $A_{1}=1/3$, $A_{2}=\sqrt{2}/3$, $c_{1}=0$, $c_{2}=c_{3}=1$.
If $c_{1}\neq 0$, the asymptotically leading polynomial in $\chi_{1}$, $\chi_{2}$ and $\Xi$ is $x^{2}+\tilde{\gamma}^{2}(3\alpha_{1}\alpha_{2})^{-2}t^{2}$. Excluding $\mathrm{Re}(\tilde{\gamma})=0$, which is $(\alpha_{1}+\alpha_{2})\mathrm{Im}(\gamma)=-1$, this polynomial is different from zero for all $x,t$. Then we obtain
$u^{\prime}_{\mu}\sim\frac{1}{2}A_{\mu}\,e^{\mathrm{i}\,\varphi_{\mu}}\Big(-1\pm(-1)^{\mu}\,\mathrm{i}\,\sqrt{3}\Big)\qquad\quad\mu=1,2\,$,
as $x\to\infty$ or $t\to\infty$. Here the $\pm$ corresponds to that in the denominator of (6.69). As a consequence, asymptotically we have $|u^{\prime}_{\mu}|=|A_{\mu}|$. Hence the components of the solution $u^{\prime}$ are both localized on the respective plane wave background, which is the characteristic property of rogue waves (interpreting $x$ and $t$ as “space” and “time” coordinates, or vice versa). The components $u_{1}^{\prime}$ and $u_{2}^{\prime}$ form pairs of (elementary) rogue waves. An example is shown in Fig. 13.
@@
Figure 13: Plots of the absolute squares of the two components of a solution from the class in Example 6.15, exhibiting rogue wave pairs. Here we chose the same data as in Fig. 12, but now $c_{1}=1$ and $c_{3}=0$.
@@
Again, we assume that $r$ is a triple root of the cubic equation (6.62), hence given by (6.68). Then $\gamma$ and $A_{\mu}$, $\mu=1,2$, are given by (6.69) and (6.70), respectively. Now the linear system takes the form
@@
where $\theta$ is given by (6.71) and $f=(f_{1},\ldots,f_{6})^{T}$ satisfies $f_{t}=Cf$ with the constant nilpotent matrix
@@
Inserting the expressions for $\chi_{1},\chi_{2}$ and $\Omega^{-1}$ in (6.61), determines a family of quasi-rational solutions of the two-component vector FL equation, which are localized on the plane wave background. These are thus rogue waves. They are of second order in the sense that here $\Gamma$ is a $2\times 2$ Jordan block, which causes the appearance of higher (than in the $n=1$ case) order polynomials of the dependent variables. $\Box$